# Louis Serrano1  Armand Kassaï Koupaï1  Thomas X Wang1 Pierre Erbacher 2Patrick Gallinari1,3 1Sorbonne Université, CNRS, ISIR, 75005 Paris, France

### v2

@@ Appendix D Additional Quantitative results
D.1 Analysis of the generation
@@
In this section, we aim to evaluate whether our pretrained model can generate new samples given the observation of a trajectory in a new environment. The key difference with previous settings is that we do not condition the transformer with tokens derived from a real initial condition. We expect the model to generate trajectories, including the initial conditions, that altogether follow the same dynamics as in the observations. Our objective is to assess three main aspects: 1) Are the generated trajectories faithful to the context example, i.e., do they follow the same dynamics as those observed in the context ? 2) How diverse are the generated trajectories—are they significantly different from each other? 3) What type of initial conditions does Zebra generate?
@@
To quantify the first aspect, we propose a straightforward methodology. We generate ground truth trajectories using the physical solver that was originally employed to create the dataset, starting with the initial conditions produced by Zebra and using the ground truth parameters of the environment (that Zebra cannot access). We then compute the $L^{2}$ distance between the Zebra-generated trajectories and those generated by the physical solver. For the second aspect, we calculate the average distance between the Zebra-generated trajectories to measure diversity. These two metrics are presented in Table 7 for both the Advection and Combined Equations. Finally, as a qualitative analysis, we perform PCA on the initial conditions generated by Zebra, and we visualize the first two components in Figure 13 for the Combined Equation case.
@@
We keep the default temperature ($\tau=1.0$) to put the focus on diversity, and for each context trajectory, we generate 10 new trajectories in parallel.
@@ Results
According to table 7, we can conclude that in this context, Zebra can faithfully generate new initial conditions and initial trajectories that respect the same physics as described in the context example. This means that our model has learned the statistical properties that relate the initial conditions with the later timestamps. The high average $L^{2}$ between generated samples indicate that the generated samples are diverse. We can visually verify this property by looking at fig. 13, noting that the generated samples cover well the distribution of the real samples.
Table 7: Fidelity and diversity - The $L^{2}$ is a proxy score for measuring the fidelity to the dynamics in the context. The average $L^{2}$ between samples quantifies the distance between each generation.
@@
(a) Analysis of the distribution of the generated initial conditions (t=0).
(b) Analysis of the distribution of the generated trajectories (t=9).
Figure 13: Qualitative analysis. We generate new initial conditions and obtain rollout trajectories with Zebra on new test environments. We then perform a PCA in the physical space to project on a low-dimensional space, at two given timestamps to check whether the distributions match.
@@ Appendix E Qualitative results
We provide visualizations of the trajectories generated with Zebra under different settings in the following figures:
@@ E.1 Advection
Figure 15: Zero-shot prediction on Advection
Figure 16: Five-shot adaptation on Advection
@@ E.2 Burgers
Figure 19: Zero-shot prediction on Burgers
Figure 20: Five-shot adaptation on burgers
@@ E.3 Heat
Figure 23: Five-shot prediction on Heat
Figure 24: Zero-shot adaptation on Heat
@@ E.5 Combined equation
Figure 30: Zero-shot prediction on Combined
Figure 31: Five-shot adaptation on Combined
@@
E.6 Vorticity
Figure 33: One-shot adaptation on Vorticity. Example 1.
Figure 34: One-shot adaptation on Vorticity. Example 2.
Figure 35: One-shot adaptation on Vorticity. Example 3.
Figure 36: Zero-shot prediction on Vorticity. Example 1.
Figure 37: Zero-shot prediction on Vorticity. Example 2.
Figure 38: Zero-shot prediction on Vorticity. Example 3.
@@ E.7 Wave 2D
Figure 42: Zero-shot prediction on Wave2d. Example 1.
Figure 43: Zero-shot prediction on Wave2d. Example 2.
Figure 44: Zero-shot prediction on Wave2d. Example 3.

### v3

@@ C.4 VQVAE
The quantizer used at the token level is a VQVAE model (Oord et al., 2017). As illustrated in Figure 10 this is an encoder-decoder architecture with an intermediate quantizer component.
The encoder spatially compresses the input function ${\bm{u}}^{t}$ by reducing its spatial resolution $H\times W$ to a lower resolution $h\times w$ while increasing the channel dimension to $d$. This is achieved through a convolutional model $\mathcal{E}_{w}$, which maps the input to a continuous latent variable $\mathbf{z}^{t}=\mathcal{E}_{w}({\bm{u}}^{t})$, where $\mathbf{z}^{t}\in\mathbb{R}^{h\times w\times d}$. The latent variables are then quantized to discrete codes $\mathbf{z}_{q}^{t}$ using a codebook $\mathcal{Z}$ of size $K=|\mathcal{Z}|$ and through the quantization step $q$. For each spatial code $\mathbf{z}^{t}_{[ij]}$, the nearest codebook entry $z_{k}$ is selected:
The training objective is:
where the first term is the Relative L2 loss commonly used in PDE modeling, and the second term is the commitment loss, ensuring encoder outputs are close to the codebook entries. The parameter $\alpha$, set to 0.25, balances the two components. Here, sg denotes the stop-gradient operation that detaches a tensor from the computational graph.
@@
Hyperparameters Advection Heat Burgers Wave b Combined Vorticity Wave 2D
@@ Appendix D Additional Quantitative results
D.1 Alternative pretrainings
We experimented with different pretraining strategies before settling on the pretraining approach proposed in Zebra. The most intuitive way to adapt the next-token prediction objective for dynamics modeling is to operate in a continuous latent space, omitting the quantization step used in Zebra and therefore using a deterministic transformer instead of a generative model. As shown in previous studies (Li et al., 2024; Agarwal et al., 2025), we obtained better reconstruction results using an autoencoder instead of a VQVAE.
However, we encountered two critical challenges with this approach. First, the training loss plateaued quickly, as illustrated in Figure 13. Compared to training a generative transformer (with the negative log-likelihood) as in Zebra, the deterministic variant trained with MSE loss exhibited instability and failed to improve over training steps.
(a) Training loss (negative log-likelihood) with Zebra
(b) Training loss (MSE) with deterministic transformer
Figure 13: Comparison of the training process between the Zebra transformer and a deterministic transformer on Advection. In both cases, the model is trained to predict the next token. Zebra utilizes a discrete vocabulary and learns a probability distribution over the next token, whereas the deterministic transformer is optimized to predict the mean using MSE loss.
Second, while the model was able to predict the next token at inference, it could not generate an entire trajectory. It performed particularly poorly in the one-shot adaptation setting. As demonstrated in Figure 14, errors accumulated quickly during inference, leading to rapid divergence from the ground truth. This ultimately resulted in poor reconstructions when feeding the predicted tokens into the decoder, as seen in Figure 15.
Figure 14: When trained with MSE for next-token prediction, inference suffers from instabilities, causing errors to grow exponentially. Here, we show the evolution over the sequence tokens of a particular channel to illustrate the phenomenon. The dashed lines show the temporal transitions between two subsequent frames, as here each frame is represented by 16 tokens.
Figure 15: The deterministic transformer, trained with the next-token objective, performs poorly in the one-shot adaptation task.
These observations highlight the importance of the generative aspect when adopting a next-token objective. For this reason, we opted for a quantized representation combined with a transformer modeling a discrete distribution, a standard approach in image and video generation, but which has never been explored for modeling physical phenomena. While alternative strategies exist (Tian et al., 2024), they involve non-trivial extensions and are left for future works.
That said, next-token prediction pretraining may not be the only viable framework for developing in-context capabilities. To explore this, we experimented with a direct next-frame prediction approach using a deterministic setup trained with relative L2 loss. This method, which we called VIT-in-context serves as a baseline for evaluating other in-context pretrainings. It is based on a video transformer operating on patches with bidirectional attention. While its training behavior was more stable, inference results remained unsatisfactory.
D.2 Uncertainty quantification
Figure 16: Uncertainty quantification with Zebra in a one-shot setting on Heat equation
@@ D.2 Uncertainty quantification
Motivated by this observation, we investigate how varying the model’s temperature parameter $\tau$ affects its predictions; specifically in the one-shot adaptation setting described in Section 4.2. By adjusting $\tau$, we aim to assess its impact on both the accuracy and variability of the predictions. We generate 10 different trajectories and use them to compute several metrics. We employ four metrics to evaluate the model’s uncertainty:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    Continuous Ranked Probability Score (CRPS, Gneiting & Raftery (2007) ) is a proper scoring rule that measures the accuracy of a probabilistic forecast by quantifying the difference between the predicted cumulative distribution function (CDF) and the empirical CDF of the observed value, with lower values indicating better calibration and sharpness.
  5. 5.
    Root Mean Squared Calibration Error (RMSCE, Hendrycks et al. (2018)) quantifies the discrepancy between predicted confidence and actual accuracy at a given confidence level.
Results
When modeling uncertainty, the model achieves a tradeoff between the quality of the mean prediction approximation and the guarantee for this prediction to be in the corresponding confidence interval. Figure 17 illustrates the trade-off between mean prediction accuracy and uncertainty calibration. At lower temperatures, we achieve the most accurate predictions, but with lower variance, i.e. with no guarantee that the target value is within the confidence interval around the predicted mean. Across most datasets, the confidence level then remains low (less than 80% for $\tau<0.25$), indicating that the true solutions are not reliably captured within the empirical confidence intervals. Conversely, increasing the temperature results in less accurate mean predictions and higher relative standard deviations, but the confidence intervals become more reliable, with levels exceeding 95% for $\tau>0.5$. Therefore, the temperature can be calibrated depending on whether the focus is on accurate point estimates or reliable uncertainty bounds.
To better calibrate this temperature, we can therefore use a proper scroring metric such as the CRPS, and we can pick the temperature parameter that has the lowest CRPS value for a given value (see Figure 17). We can see that Zebra models really well the distribution for Combined and Advection and not so much for Burgers and Heat somehow.
Finally, we examine how the model’s uncertainty evolves as additional information is provided as input. Specifically, we compare Zebra’s average error and relative uncertainty when conditioned on one example trajectory, with one or two frames as initial conditions. Table 9 reports the relative L2 loss and relative standard deviation for both scenarios. The results clearly show that including the first two frames as initial conditions reduces both the error and the relative standard deviation consistently. This indicates that, while some of the uncertainty remains aleatoric, the epistemic uncertainty is reduced as more input information becomes available.
Table 9: Uncertainty quantification in the one-shot setting. Conditioning from a trajectory example and 1 frame or 2 frames as initial conditions. Metrics include relative $L^{2}$ loss (average accuracy) and relative standard deviation (average spread around the average prediction). The temperature is fixed at 0.1.
Advection Heat Burgers Wave b Combined
Rel. $L^{2}$ 1 frame 0.006 0.156 0.115 0.154 0.008
Rel. $L^{2}$ 2 frames 0.004 0.047 0.052 0.075 0.005
Rel. Std. 1 frame 0.003 0.062 0.048 0.074 0.005
Rel. Std. 2 frames 0.002 0.019 0.018 0.040 0.003
Figure 17: Uncertainty quantification with Zebra. The main parameter of this study is the temperature (x-axis). We then look from left to right at (1) The rollout loss, i.e. the relative $L^{2}$ loss between the predictions and the ground truth; (2) The relative standard deviation to quantify the spread around the mean; (3) The confidence level, that measures the frequency of observations that lie within the empirical confidence interval. (4) The CRPS that measures the quality of the uncertainty, can be used to pick the temperature with the most calibrated uncertaity.
Comparison with ViT-in-Context
We quantitatively evaluate the uncertainty using CRPS and RMSCE in the one-shot prediction task, as reported in LABEL:tab:crps-rmsce-comparison, comparing Zebra with ViT-in-Context. Since ViT-in-Context is a deterministic model, we introduce stochasticity by: (1) adding Gaussian noise with a small standard deviation ($\sigma=0.1$) to the input, thereby perturbing the input; and (2) enabling stochastic depth at inference time by keeping the DropPath mechanism (Huang et al., 2016). Across all settings, Zebra consistently outperforms these simpler baselines, often by an order of magnitude in both CRPS and RMSCE.
Table 10: CRPS and RMSCE results across PDE benchmarks. Lower is better.
Metric Model Advection Heat Burgers Wave b Combined
CRPS ViT + noise 0.0705 0.176 0.227 0.093 0.098
ViT Dropout 0.0363 0.213 0.196 0.024 0.024
Zebra 0.0026 0.043 0.020 0.0129 0.0018
RMSCE ViT + noise 0.132 0.241 0.265 0.249 0.045
ViT Dropout 0.386 0.547 0.529 0.340 0.064
Zebra 0.074 0.055 0.048 0.124 0.074
Uncertainty over time
Finally, we illustrate in LABEL:tab:crps-rmsce-time how CRPS and RMSCE evolve over successive autoregressive steps in the predicted trajectory for Zebra on the Burgers equation. As expected, CRPS increases over time due to error accumulation during rollout. In contrast, RMSCE remains stable, indicating that the model maintains consistent calibration throughout the prediction.
Table 11: CRPS and RMSCE over time for Zebra on the Burgers dataset.
Metric $t=1$ $t=2$ $t=3$ $t=4$ $t=5$ $t=6$ $t=7$ $t=8$ $t=9$
CRPS 0.0057 0.0106 0.0148 0.0184 0.0216 0.0244 0.0271 0.0296 0.0320
RMSCE 0.0511 0.0513 0.0505 0.0487 0.0483 0.0489 0.0467 0.0457 0.0457
D.3 Analysis of the generation
Zebra is capable of generating completely novel trajectories for new environments, including the initial conditions. An example of a generated trajectory for Vorticity 2D is shown in Figure 18, where the top row shows the context trajectory used to guide the generation, and the bottom row displays the model’s generated trajectory, including the initial condition. In this section, we evaluate on Combined and Advection the quality of the generated trajectories with Zebra.
Figure 18: Unconditional generation on Vorticity 2D. The top-row is the example used to guide the generation, and the bottom-row is the generated example. The model also generates the initial condition.
Setting
We evaluate whether our pretrained model can generate new samples conditioned on a trajectory observed in a previously unseen test environment. Unlike previous settings, the transformer is not explicitly conditioned on tokens derived from a real initial condition. Instead, we expect it to generate trajectories, including their initial conditions, that follow the same dynamics as the observed context.
Our evaluation focuses on four key aspects. First, we assess whether the generated trajectories faithfully follow the dynamics of the context. Second, we analyze the diversity of the generated trajectories to determine if they are significantly different from one another. Third, we compare the generated samples with those produced by numerical solvers to evaluate whether their distributions align. Finally, we examine the types of initial conditions generated by Zebra.
Metrics
To assess fidelity, we generate ground truth trajectories using the physical solver originally used to construct the dataset. These simulations start from the initial conditions generated by Zebra, using ground truth environment parameters that Zebra itself does not have access to. We then compute the $L^{2}$ distance between the generated trajectories and those obtained from the physical solver.
For diversity, we measure the average pairwise $L^{2}$ distance between different trajectories generated by Zebra. The results for both fidelity and diversity are reported in Table 12 for the Advection and Combined Equations.
To compare the distribution of generated trajectories with that of numerical solvers, we compute the Wasserstein distance using the Sinkhorn algorithm (Cuturi, 2013). As baselines, we compare against two references. First, we compute the Wasserstein distance with purely random samples drawn from Gaussian noise, providing an upper bound on the problem. Second, we measure the Wasserstein distance between two independent sets of numerical solver trajectories (the validation and test sets), allowing us to quantify the variability inherent in the dataset itself.
Finally, as a qualitative analysis, we perform a principal component analysis on the trajectories generated by Zebra and visualize the first two principal components in Figure 19 for Combined.
@@ D.3 Analysis of the generation
We use a default temperature of $\tau=1.0$. For each context trajectory, we sample 10 new trajectories in parallel.
@@ Results
Table 12 shows that Zebra generates new initial conditions and trajectories that respect the same physical laws as the given context. The model seems to have learned the statistical relationships between initial conditions and later timestamps. The high average $L^{2}$ distance between samples indicates that the generated trajectories are diverse. This can also be observed in Figure 19, where the generated samples effectively cover the distribution of real samples.
Table 12: Fidelity and diversity metrics. The $L^{2}$ distance measures fidelity to the context dynamics, while the average $L^{2}$ quantifies sample diversity.
@@
To further assess distribution alignment, we compute the Wasserstein distance between the generated trajectories and those obtained with numerical solvers. The results in Table 13 indicate that Zebra achieves lower Wasserstein distances than Gaussian noise but remains slightly above the cross-distribution baseline (which compares the validation and test distributions), suggesting reasonable alignment with the true data distribution.
Table 13: Comparison of distributions using the Wasserstein distance between Zebra-generated trajectories and numerical solver samples.
Distance Metric Advection Combined Equation
Gaussian noise vs. real data 18.22 16.15
Validation data vs. test data 5.11 1.87
Zebra-generated data vs. real data 5.57 2.21
(a) Distribution of generated initial conditions ($t=0$).
(b) Distribution of generated trajectories at ($t=9$).
Figure 19: Qualitative analysis of generated trajectories. Zebra generates new initial conditions and trajectories for unseen test environments. PCA projections visualize both generated and true trajectories in a lower-dimensional space at $t=0$ and $t=9$.
D.4 Dataset scaling analysis
We investigate how the one-shot error on the test set evolves as we vary the size of the training dataset. To this end, we train the auto-regressive transformer on datasets containing 10, 100, 1000, and 12,000 trajectories and evaluate Zebra’s generations on the test set, starting with two frames as inputs. The training time is proportional to the dataset size: for example, the number of training steps for 1,000 trajectories is 10 times the number of steps for 100 trajectories. The results are presented in Figure 20.
First, we observe that Zebra requires a substantial amount of data to generalize effectively to different parameter values, even within the training distribution. This aligns with findings in the literature that transformers, especially auto-regressive transformers, excel at scaling —performing well on very large datasets and for larger model architectures. However, for smaller datasets, this approach may not be the most efficient. We believe that Zebra’s potential resides when applied to large amounts of data, making it an ideal candidate for scenarios involving large-scale training.
Second, for the Combined equation, we notice that performance plateaus between 100 and 1,000 trajectories. This may be due to insufficient training or a lack of diverse examples, as the Combined equation is more challenging compared to the Advection equation, whose performance follows a more log-linear trend. This suggests that additional data or targeted training strategies might be needed to achieve better generalization for more complex equations.
Figure 20: Dataset scaling analysis. One-shot prediction error on the test set vs. the training dataset size.
D.5 Inference time comparison
Table 14 compares the inference time for one-shot adaptation across different methods when predicting a single trajectory given a context trajectory and an initial condition. For Zebra, the inference process, which includes encoding, auto-regressive prediction, and decoding, is much faster in 1D and slightly faster in 2D. With Zebra, the bottleneck at inference is the autoregressive generation of tokens, which speed is about 128 tokens per second on a V100 for 2D and an RTX for 1D. The decoding is fast and can be done in parallel for the trajectory in one forward pass. In contrast, for CODA and CAPE, the majority of the inference time is spent on adaptation and gradient-based steps. Here the times were reported with 100 gradient steps, note that we used 250 for the rest of the experiments. We believe Zebra’s inference time could be further optimized by (1) improving the optimization code and leveraging specialized hardware such as H100 (for flash attention) and LPUs (which show significant speed-ups against GPUs), and (2) increasing the number of tokens sampled per step (as in e.g. next-scale prediction (Tian et al., 2024)).
However, we have shown that it is possible to greatly accelerate inference by employing a deterministic neural surrogate on top of Zebra, which acts as a dynamic encoder. This framework is order of magnitudes faster than gradient-based adaptation methods.
Table 14: Inference times for one-shot adaptation. Average time in seconds to predict a single trajectory given a context trajectory and an initial condition. Times include adaptation and forecast for CODA and CAPE, while it includes encoding, auto-regressive prediction and decoding for Zebra.
Advection Vorticity 2D
CAPE 18s 23s
CODA 31s 28s
Zebra 3s 21s
Zebra + UNet 0.10s 0.14s
D.6 Influence of the codebook size
The codebook size $K$ is a crucial hyperparameter. It directly affects the quality of the reconstructions, since a larger codebook can improve the reconstructions quality. However, it also impacts the dynamics modeling stage: the smaller the codebook, the easier it is for the transformer to learn the statistical correlations between similar trajectories. To have a sense of this trade-off, we report the relative reconstruction errors and the one-shot prediction errors in Table 15. The reconstruction error decreases when the codebook size increases. However, the one-shot prediction error decreases from 32 to 64 codes but then gradually increases from 64 to 512. We can see that it follows a U-curve in Figure 21. This phenomenon was observed in a different context in (Cole et al., 2024).
Table 15: Influence of the codebook size. Reconstruction error and one-shot prediction error on Burgers for different codebook sizes. Errors in relative L2.
Codebook Size Reconstruction One-shot Prediction
32 0.0087 0.116
64 0.0043 0.097
128 0.0024 0.124
256 0.0019 0.163
512 0.0015 1.093
Figure 21: One-shot accuracy vs codebook size. One-shot prediction error on the test set for various codebook sizes. Error in relative L2.
D.7 Reconstruction errors
We report the accuracy of the reconstructions from our encoder-decoder for different datasets in LABEL:tab:vqvae-reconstruction-scores. Here, no dynamics is involved, we simply evaluate the quality of the encoding and of the decoding. On 1D and 2D datasets, the decoding errors are respectively of 0.1 % and 1% on the test set.
Table 16: Reconstruction errors. Test relative L2 loss between reconstructions from Zebra’s VQVAE and the ground truths.
Advection Heat Burgers Wave b Combined Wave 2D Vorticity 2D
VQVAE of Zebra 0.0003 0.0019 0.0016 0.0011 0.0022 0.010 0.017
To assess how faithful signal encoding and deconding influences prediction accuracy, we evaluate Zebra in an out-of-distribution setting on the Vorticity 2D dataset across several viscosity ranges. As viscosity decreases and departs from the training distribution, both the context and target trajectories exhibit increasingly high-frequency components which were not seen by the VQVAE. We report reconstruction scores alongside one-shot prediction errors in Table 17.
Table 17: Impact of encoding and decoding quality on prediction accuracy. Reconstruction and one-shot prediction error with Zebra. Metric is relative L2.
Viscosity Range Type Reconstruction One-shot Prediction
$[10^{-3},10^{-2}]$ In-distribution 0.02 0.12
$[5\times 10^{-4},10^{-3}]$ Close OoD 0.13 0.24
$[10^{-5},10^{-4}]$ Far OoD 0.22 0.32
As the viscosity decreases and deviates from the training distribution, the reconstruction error increases, which is expected since the VQVAE was not exposed to these regimes. However, the one-shot prediction error remains stable. This suggests that the encoder continues to capture the essential characteristics of the context dynamics in this one-shot OoD setting, enabling the transformer to leverage these sequence indices for accurate forecasting of the low-frequency components.
D.8 Influence of the number of context examples
While all models were evaluated on a one-shot adaptation task, Zebra was trained with up to 5 context examples as input. This allows us to analyze the impact of context size on prediction accuracy. As shown in Table 18, performance saturates after approximately 3 in-context examples.
Table 18: Effect of number of context example. Few-shot prediction error (relative L2) on 1D datasets as a function of the number of context examples.
# Examples Advection Heat Burgers Combined
1 0.0074 0.1563 0.1150 0.0095
2 0.0077 0.1310 0.1020 0.0074
3 0.0072 0.1272 0.1000 0.0078
4 0.0071 0.1272 0.0990 0.0075
5 0.0071 0.1310 0.1000 0.0073
@@ Appendix E Qualitative results
We provide visualizations of the trajectories generated with Zebra under different settings in the following figures. One-shot prediction: Figure 22, Figure 24, Figure 26, Figure 28, Figure 30, Figure 32, Figure 38. Uncertainty quantification: Figure 23, Figure 25, Figure 27, Figure 29, Figure 27.
@@ E.5 Combined equation
E.6 Vorticity 2D
Figure 32: One-shot adaptation on Vorticity 2D. Example 1.
Figure 33: One-shot adaptation on Vorticity 2D. Example 2.
Figure 34: One-shot adaptation on Vorticity 2D. Example 3.
E.6.1 Out-of-distribution
Figure 35: One-shot OoD adaptation on Vorticity 2D. Example 1.
Figure 36: One-shot OoD adaptation on Vorticity 2D. Example 2.
Figure 37: One-shot OoD adaptation on Vorticity 2D. Example 3.
@@ E.7 Wave 2D
Figure 40: One-shot adaptation on Wave 2D. Example 3.